Calculate the degree of ionisation and the pH of a 0.01 M aqueous solution of formic acid at 298K. [Given, Ka (HCOOH) = 1.7  10−4at 298K]. 


1
Expert's answer
2022-03-19T08:52:42-0400

"HCOOH \leftrightarrows H^+ +HCOO^-"


When equilibrium is reached the concentrations of each form will be:


c(HCOOH) = 0.01 - x M,

c(H+) = x M,

c(HCOO-) = x M, where x is the dissociated portion of the acid


"Ka(HCOOH)=1.7*10^{-4}=\dfrac{c(H^+)c(HCOO^-)}{c(HCOOH)}=\dfrac{x^2}{0.01-x}"


This equation gives "x =0.0012" M


Then the degree of ionisation: "\alpha=\dfrac{c(H^+)}{c(HCOOH)}=\dfrac{0.0012}{0.01}=0.12"


"pH=-log_{10}c(H^+)= -log_{10}0.0012=2.92"